# Incorporating Gradients to Rules: Towards Lightweight, Adaptive Provenance-based Intrusion Detection

### v1

@@ IV-A System Overview
As shown in Fig. 3, Captain consists of two major parts: the detection module and the learning module. The detection module takes the audit logs as input and produces detection results, while the learning module leverages the false alarms produced to fine-tune the detection module during the training phase.
@@ IV-B Adaptive Detection Module
As shown in Fig. 3, the detection module in Captain comprises three major components: tag initialization, tag propagation, and alarm generation. The tags contain the security states and contextual information of system entities. System entities are assigned initial tags when they appear in the system. As additional system events happen, tags propagate among system entities accordingly. At the same time, Captain determines whether to generate alarms by comparing the tags and the malicious threshold.
Tag Design: As shown in Table II, Captain uses two types of tags: data tags and code tags, which correspond to the data flow and control flow, respectively. Data tags exist on all nodes in the provenance graph, representing the maliciousness of the data. A data tag is represented as a numerical vector <c, i>, where c denotes the confidentiality score of the data, and i denotes the integrity score of the data. In contrast, code tags only exist on process nodes, indicating the maliciousness of the code. A code tag holds a numerical value <p>, where <p> denotes the integrity score of the code.
All scores (c, i, and p) are real numbers ranging from 0 to 1. A low confidentiality score is associated with highly sensitive data, such as passwords, while a high confidentiality score corresponds to public data. A low integrity score suggests that the data/code might be compromised and, therefore, is considered untrusted, while a high integrity score implies that the data/code is trusted.
@@
Moreover, we introduce three adaptive parameters:
@@
Tag Initialization: Proper tag initialization rules play a crucial role since they determine the initial security states of system entities. For instance, if a socket associated with a benign IP address is assigned low integrity, it could lead to many false alarms. However, it is relatively less discussed in many previous works [19, 18]. They usually assign the initial tags based on domain knowledge without further adjustments. Although Morse admits tag initialization policies could be learned from the previous tracing data, it did not discuss it in detail or propose a practical methodology.
@@
For non-process nodes, we define the adaptive parameter $A$ at the node level to facilitate fine-grained tag initialization rules. Specifically, for any node $n_{i}\in N$ in the provenance graph, $A$ stores its initial data tag with respect to its node feature. For example, the node features for a socket are its IP address. Then, all sockets with the same IP address will be assigned the same initial data tag.
The adaptive parameter $A$ allows Captain to learn the fine-grained tag initialization rules at the node level. Captain’s learning module refines $A$ through the training process, which will be elaborated upon in §IV-C. The training process starts from the default value $A_{0}$. Since we train Captain on the benign data, we set $A_{0}$ conservatively. In other words, we want to ensure all potential attacks can be captured with our initial parameter $A_{0}$. For example, we assign all IP addresses an initial integrity of 0 so we do not miss any attackers entering through network communications.
@@
In Captain, tag propagation will pass and accumulate malicious intentions. Unfortunately, such a tag propagation mechanism can suffer from the dependence explosion issue [19] and lead to bad detection performance. For example, as shown in III, (bash, write, /home/user/.bash_history) is a common event that appears in many propagation chains of false alarms, which means this event causes large-scale maliciousness propagation on the provenance graph. However, our investigation shows this is a commonly seen benign activity.
To mitigate such a problem, we introduce another adaptive parameter $G$, to regulate the propagation rate. Specifically, for any edge $e_{i}\in E$ identified with (src_node_feature, event_type, dest_node_feature) in the provenance graph, $G$ stores its propagation rate parameter $g_{i}\in[0,1]$. Given $g_{i}$, the source node $src$, and the destination node $dest$, we update $tag_{dest}$ as follows during tag propagation.
$tag_{dest}^{new}=g_{i}\cdot tag_{rule}+(1-g_{i})\cdot tag_{dest}$ (1)
Where $tag_{rule}$ is the tag value given by the propagation rules defined in Table XI. Usually, it equals to $tag_{src}$. If $g_{i}$ is close to 0, the corresponding tag propagation will lead to small changes in the tag values and vice versa. In this way, $G$ allows Captain to fine-tune the propagation rules at the edge level.
Captain refines $G$ through the training process in the learning module, which will be elaborated upon in §IV-C. We also conservatively establish the default values of $G$ ($G_{0}$) to ensure all true positive alarms can be captured before the training starts. Particularly, we set $g_{i}=1$ for all edge $e_{i}\in E$ identified with (src_node_feature, event_type, dest_node_feature) to make all propagation fully effective.
@@
We introduce an adaptive parameter $T$ to enable fine-grained adjustment of the alarm threshold. Specifically, for any edge $e_{i}\in E$ identified with (src_node_feature, event_type, dest_node_feature) in the provenance graph, $T$ stores its alarm threshold $thr_{i}\in[0,1]$. For an edge $e_{i}$, the detection function is defined as follows.
$f(e_{i})=tag-thr_{i}=\begin{cases}malicious,&\text{if }f(e_{i})<0\\ benign,&\text{otherwise}\\ \end{cases}$ (2)
where $tag$ is the relevant tag on the node of interest on edge $e_{i}$. A low threshold prevents the system from generating the alarms, while a high threshold encourages the system to generate the alarms. In this way, $T$ allows Captain to adjust the alarm threshold at the edge level.
Like $A$ and $G$, Captain’s learning module refines $T$ can be refined through the training process, which will be elaborated upon in §IV-C. Before the training, we set the default thresholds ($T_{0}$) neutrally. In other words, we set $thr_{i}=0.5$ for every edge $e_{i}$ so that we do not encourage or suppress all alarms.
IV-C Configuration Learning Module
With the three adaptive parameters ($A$, $G$, and $T$) we defined in the previous section, the problem of adaptative configuration for the detection system is then converted to the optimization problem of the multi-variable function aiming to find the optimal values for $A$, $G$, and $T$. In this section, we present our efforts to solve this optimization problem with the gradient descent algorithm, including defining the objective function and constraints, calculating gradients, and searching for the optimal values.
@@ IV-C1 Loss Function
The false alarm term mainly focuses on penalizing erroneously triggered false alarms. More specifically, an event $e$ triggers a false alarm means $f(e)$ should be greater than 0 but it does not. Since $f(e)\in[-1,1]$, we can use the Mean Squared Error $(y_{e}-f(e))^{2}$ as the loss function for all events that trigger a false alarm. $y_{e}$ is set as $1$ for those benign events, and $-1$ for the malicious events. Because we only use benign data for training, all $y_{e}$ should be $1$. Please also note that we do not calculate the loss for the correctly classified events, i.e., we do not make $f(e)$ close to $1$ if $f(e)$ is already greater than 00. This is because the events that do not cause alarms are significantly more than the alarm-triggering events, considering them, therefore, is inefficient and would make the detection system insensible to the malicious events. Thus, the false alarm term in the loss function can be formalized as
@@
To capture the potential malicious events as much as possible, the adaptive parameters are configured conservatively. During training, we only want to make small adjustments to a small portion of the parameters according to the benign data to reduce the false alarms. For the rest parameters, we would like to keep them as conservative as possible so that the sensitivity to the maliciousness is not compromised.
In the area of machine learning, One-class classification (OCC) algorithms are proposed to deal with the situation where only one class of samples is available in the training set [20]. Inspired by [40, 41], we add a regularizer term in the loss function to avoid the parameters becoming too lenient when the malicious training sample is absent. The essential thought is to make the adaptive parameters as close to the default values ($A_{0},G_{0},T_{0}$) as possible. By minimizing the $l_{2}$ distance between the adaptive parameters and the default values, we protect the detection capability during training. Therefore, the detection loss can be formalized as
$\small Loss=\sum_{e\in E}\mathcal{L}(e)+\alpha||A-A_{0}||_{2}+\gamma||G-G_{0}||_{2}+\tau||T-T_{0}||_{2}$ (4)
where, $\alpha$, $\gamma$, and $\tau$ are the regularizer coefficients.
@@ IV-C2 Differentiable Detection Framework
One of the fundamental steps of using gradient descent algorithms is calculating the gradient of each variable. To adjust the adaptive parameters based on the loss function, we need to implement a differentiable rule-based PIDS, which allows us to keep the gradients of the adaptive parameters with respect to the $Loss$.
@@
$\frac{\partial Loss}{\partial a_{n}}=\frac{\partial\mathcal{L}}{\partial f}\cdot\frac{\partial f}{\partial a_{n}}+\alpha$ (5)
The equations for $g_{e}$ and $thr_{e}$ hold similarly to Eq. 5.
As mentioned before, the most challenging part of building an adaptative rule-based PIDS is to compute the gradients, i.e., $\partial f/\partial a_{n}$, $\partial f/\partial g_{e}$, and $\partial f/\partial thr_{e}$. This is because the rule-based detection process is usually modeled as the branch selection based on rules, rather than a differentiable function. To the best of our knowledge, no previous work has formalized or recorded the gradients of parameters in a tag-based PIDS. In Captain, we model the detection process as a differentiable function and calculate the gradients of each adaptive parameter defined in § IV-B. When the tags are propagated, the corresponding gradients are updated and recorded. In so doing, we associate the adaptive parameters with the loss, making it possible to perform the gradient descent algorithm to find the optimal parameters that minimize the loss.
As per Eq. 2, $\partial f/\partial thr_{e}=-1$. This means every time when we want to increase the value of $f$ (the direction of benign), we have to decrease the value of $thr_{e}$, and vice versa. This aligns with our intuition because in our framework, a lower tag value means more maliciousness. Therefore, if we want to be more lenient in the detection, we should set a lower threshold.
The formalization and calculation of $\partial f/\partial a_{n}$ and $\partial f/\partial g_{e}$ are more complicated. According to Eq. 2, calculating the gradients of $f$ is essentially computing the gradients with respect to the variable $tag$ for each node.
We start from the gradients of $a_{n}$. According to the propagation policies in table, $tag_{rule}$ is either $min(tag_{src},tag_{dest})$ or a constant $c$. If $tag_{rule}=c$, then
@@
Please note that if $tag_{rule}=tag_{dest}$, i.e. we do not need to update the tags of the $dest$ node, the gradients of $dest$ are unchanged as well.
@@
$\small\frac{\partial tag_{n^{\prime}}}{\partial a_{n}}=\begin{cases}1,&\text{if }n^{\prime}=n\\ 0,&\text{if }n^{\prime}\neq n\end{cases}$ (8)
Then we focus on the gradient of $g_{e}$, which is more complex because it exists in the equation explicitly. Therefore, the equations of calculating the gradients are slightly different. Here we use $M$ to represent $min(tag_{src},tag_{dest})$ for the succinctness.
$\frac{\partial tag_{dest}^{new}}{\partial g_{e^{\prime}}}$ =$=\begin{cases}g_{e}\frac{\partial M}{\partial g_{e^{\prime}}}+(1-g_{e})\frac{\partial tag_{dest}}{\partial g_{e^{\prime}}},e\neq e^{\prime}\\ g_{e}\frac{\partial M}{\partial g_{e}}+(1-g_{e})\frac{\partial tag_{dest}}{\partial g_{e}}+M-tag_{dest},e=e^{\prime}\end{cases}$ (9)
@@
We use Eq. 10 to initialize the gradient of $g_{e}$ and update them according to Eq. 9. Besides, similar to the case of $a_{n}$, when the tag is not updated, i.e. $M=tag_{dest}$, we do not have to update the gradients on the related nodes.
In summary, to employ the gradient descent algorithm, we calculate the gradients of the adaptive parameters with respect to the loss. We first calculate the initial gradient for each node. Afterward, the gradients are updated according to our equations when the tags are propagated.
@@ IV-C3 Training and Testing
In the previous section, we create a differentiable tag propagation framework to calculate the gradients of the adaptive parameters with respect to the loss. We can now utilize the gradient descent algorithm to optimize the adaptive parameters. In machine learning, this process is commonly referred to as “training”. The learned parameters are stored as a customized configuration, which is then used to set up our system prior to testing.
As shown in Fig. 3, an epoch in training comprises the forward and backward propagation. Before we start learning, we configure all adaptive parameters, $A$, $G$, and $T$, as the default settings $A_{0}$, $G_{0}$, and $T_{0}$. Instead of random initialization[8], we want to use the most conservative setting as the starting point to keep the sensitivity to the maliciousness during training.
When Captain processes the events, tags are propagated among the graph, updating the gradients according to Eq. 7 and Eq. 9, and generate the detection results. Next, we calculate the loss according to the detection results and back-propagate the gradients of loss according to Eq. 5. Finally, we update the parameters in the opposite direction of the gradient as follows:
@@
where $p$ is the adjustable parameters (which could be $a_{n}$, $g_{e}$, or $thr_{e}$), $p_{new}$ is the new parameters, $p_{old}$ is the old one, $\frac{\partial Loss}{\partial p_{old}}$ is the gradient, and $l$ is the learning rate. We repeat the training process when the maximum epoch is reached or the changes in the result become sufficiently small.
Before testing, we configure the parameters according to what we learned in the training stage. Then, Captain can process audit data and conduct detection as introduced in § IV-A.

### v2

@@ IV-A System Overview
As shown in Fig. 3, Captain consists of two major parts: the detection module and the learning module. The detection module takes the audit logs as input and produces detection results, while the learning module leverages the false alarms during the training phase to fine-tune the detection module.
@@ IV-B Adaptive Detection Module
The detection module in Captain comprises three major components: tag initialization, tag propagation, and alarm generation. System entities are assigned initial tags when they appear in the system. As system events happen, tags get updated among system entities, which we refer to as “tag propagation”. At the same time, Captain determines whether to generate alarms based on the pre-defined rules.
Tag Design: Inspired by previous provenance-based methodology based on data flow and control flow [43, 44, 21, 22], Captain designs two types of tags: data tags and code tags. As shown in Table II, data tags exist on all nodes in the provenance graph. A data tag is represented as a numerical vector <c, i>, where c denotes the confidentiality score of the data, and i denotes the integrity score of the data. In contrast, code tags only exist on process nodes, denoted by <p> to indicate the integrity of the code. All scores (c, i, and p) are real numbers ranging from 0 to 1. A low integrity score suggests that the data/code might be compromised and, therefore, is considered untrusted. A low confidentiality score is counterintuitively associated with highly sensitive data (to make the scoring scheme consistent), such as passwords.
@@
The tag system used by Captain is inspired by previous rule-based PIDS [22, 21]. The focus of this paper is not on the design of the tag system. In essence, any new type of tag can be integrated with Captain as long as it satisfies the following conditions: 1) The tags are numerical; 2) The tags are updated through arithmetic calculation; 3) Alarms are triggered based on the tag values. For example, we can define an additional “exploitability” tag for all web service nodes and propose the corresponding propagation rules and alarm-triggering rules. We leave tag designing as a practical problem in specific scenarios and use the tags mentioned above to show the improvement purely brought by Captain’s learning module.
We then introduce three adaptive parameters, which are the core of the fine-grained rule-based PIDS:
@@
Tag Initialization: Proper tag initialization rules play a crucial role since they determine the initial states of system entities. For instance, if a socket associated with a benign IP address is assigned low integrity, it could lead to many false alarms. However, it is relatively less discussed in previous works [22, 21]. They usually assign the initial tags based on domain knowledge without further adjustments. Although Morse admits tag initialization policies could be learned from the previous tracing data, it did not discuss it in detail or propose a practical methodology.
@@
Captain’s learning module refines $A$ through the training process, which will be elaborated upon in §IV-C. The training process starts from the default value $A_{0}$. Since we train Captain on the benign data, we set $A_{0}$ conservatively. In this paper, “conservative” means the assumption of the worst-case. Under conservative settings, we aim to capture true attacks as much as possible without considering excessive false alarms. Rule-based PIDS such as Morse has to use the conservative setting to capture all potential attacks [22], while Captain only uses the conservative setting as the initial state for training on benign data. For example, we assign all IP addresses an initial integrity of 0 so we do not miss any attackers entering through network communications.
@@
Unfortunately, such a tag propagation mechanism can suffer from the dependence explosion issue [22], leading to excessive alarms. For example, as shown in III, (bash, write, /home/user/.bash_history) is a common event that appears in many propagation chains of false alarms, which means this event causes large-scale maliciousness propagation on the provenance graph. However, our investigation shows this is a commonly seen benign activity.
To mitigate such a problem, we introduce another adaptive parameter $G$, to regulate the propagation rate. Specifically, for any edge $e\in E$ identified with (src_node_feature, event_type, dest_node_feature) in the provenance graph, $G$ stores its propagation rate parameter $g_{e}\in[0,1]$. Given $g_{e}$, the source node $src$, and the destination node $dest$, we update $tag_{dest}$ as follows during tag propagation.
$tag_{dest}^{new}=g_{e}\cdot tag_{rule}+(1-g_{e})\cdot tag_{dest}$ (1)
Where $tag_{rule}$ is the tag value given by the propagation rules defined in Table VIII. Usually, it equals to $tag_{src}$. If $g_{e}$ is close to 0, the corresponding tag propagation will lead to small changes in the tag values and vice versa. In this way, $G$ allows Captain to fine-tune the propagation rules at the edge level.
Captain refines $G$ through the training process in the learning module, which will be elaborated upon in §IV-C. We also conservatively establish the default values of $G$ ($G_{0}$) to ensure all true alarms can be captured before the training starts. Particularly, we set $g_{e}=1$ for all edges $e\in E$ to make all propagation fully effective.
@@
We introduce an adaptive parameter $T$ to enable fine-grained adjustment of the alarm threshold. Specifically, for any edge $e\in E$ identified with (src_node_feature, event_type, dest_node_feature), $T$ stores its alarm threshold $thr_{e}\in[0,1]$. For an edge $e$, the detection function $f(e)$ is defined as follows.
$f(e)=tag-thr_{e}=\begin{cases}malicious,&\text{if }f(e)<0\\ benign,&\text{otherwise}\\ \end{cases}$ (2)
where $tag$ is the relevant tag on the node of interest on edge $e_{i}$. A low threshold prevents the system from generating the alarms, while a high threshold encourages the system to generate the alarms. In this way, $T$ allows Captain to adjust the alarm-triggering rules at the edge level.
Like $A$ and $G$, Captain’s learning module refines $T$ can be refined through the training process (§IV-C). Before the training, we set the default thresholds ($T_{0}$) neutrally. In other words, we set $thr_{i}=0.5$ for every edge $e_{i}$ so that we do not encourage or suppress all alarms.
Please note that Captain’s methodology is independent of specific rules or tags, making it broadly applicable to various rule-based PIDS, especially those using taint-analysis methods like Sleuth [21], Holmes [23], and Conan [24].
IV-C Learning Module
With the three adaptive parameters ($A$, $G$, and $T$) we defined in the previous section, the problem of adaptative configuration for the detector is then converted to the optimization problem of the multi-variable function aiming to find the optimal values for $A$, $G$, and $T$. In this section, we present our efforts to solve this optimization problem with the gradient descent algorithm, including defining the objective function and constraints, calculating gradients, and searching for the optimal values.
@@ IV-C1 Loss Function
The false alarm term mainly focuses on penalizing erroneously triggered alarms. Specifically, an event $e$ triggers a false alarm means $f(e)$ should be greater than 0 but it does not. Since $f(e)\in[-1,1]$, we can use the Mean Squared Error $(y_{e}-f(e))^{2}$ as the loss function for all events that trigger a false alarm. $y_{e}$ is set as $1$ for those benign events, and $-1$ for the malicious events. Because we only use benign data for training, all $y_{e}$ should be $1$. Please also note that we do not calculate the loss for the correctly classified events, i.e., we do not make $f(e)$ close to $1$ if $f(e)$ is already greater than 00. This is because the events that do not cause alarms are significantly more than the alarm-triggering events, considering them, therefore, is inefficient and would make the detection system insensible to the malicious events. Thus, the false alarm term in the loss function can be formalized as
@@
To capture all potential malicious events, the adaptive parameters are configured conservatively before training. During training, we only want to make small adjustments to a small portion of the parameters according to the false alarms on training data. For the rest parameters, we would like to keep them as conservative as possible so that the sensitivity to the maliciousness is not compromised. In the area of machine learning, One-class classification (OCC) algorithms are proposed to deal with the situation where only one class of samples is available in the training set [45]. Inspired by [46, 47], we add a regularizer term in the loss function to avoid the parameters becoming too lenient when the malicious training sample is absent. The essential thought is to make the adaptive parameters as close to the default values ($A_{0},G_{0},T_{0}$) as possible. By minimizing the $l_{2}$ distance between the adaptive parameters and the default values, we protect the detection capability during training. Therefore, the detection loss can be formalized as
$Loss=\sum_{e\in E}\mathcal{L}(e)+\alpha||A-A_{0}||_{2}+\gamma||G-G_{0}||_{2}+\tau||T-T_{0}||_{2}$ (4)
where, $\alpha$, $\gamma$, and $\tau$ are the regularizer coefficients. In the evaluation section, we discuss the effect of the regularizer coefficients (§VI-D2) and methods used to fine-tune them (§VI-A2).
@@ IV-C2 Differentiable Detection Framework
One of the fundamental steps of using gradient descent algorithms is calculating the gradient of each variable. To adjust the adaptive parameters based on the loss function, we need to design a differentiable detection framework, which allows us to keep the gradients of the adaptive parameters with respect to $Loss$.
@@
$\frac{\partial Loss}{\partial a_{n}}=\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial a_{n}}+\alpha\cdot(a_{n}-a_{0})$ (5)
The equations for $g_{e}$ and $thr_{e}$ (presented in Appendix §C) hold similarly to Eq. 5. The most challenging part of building an adaptative rule-based PIDS is to compute the gradients, i.e., $\partial f/\partial a_{n}$, $\partial f/\partial g_{e}$, and $\partial f/\partial thr_{e}$. This is because the rule-based detection process is usually modeled as the branch selection based on rules, rather than a differentiable function. To the best of our knowledge, no previous work has formalized or recorded the gradients of parameters in a rule-based PIDS. In Captain, we model the detection process as a differentiable function and calculate the gradients of each adaptive parameter. When the tags are propagated, the corresponding gradients are updated and recorded. In so doing, we associate the adaptive parameters with the loss, making it possible to perform the gradient descent algorithm to find the optimal parameters that minimize the loss.
As per Eq. 2, $\partial f/\partial thr_{e}=-1$. This means every time when we want to increase the value of $f$ (the direction of benign), we have to decrease the value of $thr_{e}$, and vice versa. This aligns with our intuition because a lower tag value means more maliciousness in our system. Therefore, if we want to be more lenient in the detection, we should set a lower threshold.
The formalization and calculation of $\partial f/\partial a_{n}$ and $\partial f/\partial g_{e}$ are more complicated. According to Eq. 2, calculating the gradients of the detection function $f$ is essentially computing the gradients with respect to the variable $tag$ for each node.
We start from the gradients of $a_{n}$. According to the propagation policies shown in Table VIII, the updated tag value $tag_{rule}$ is either the lower tag values of involved nodes $min(tag_{src},tag_{dest})$ or a constant value $c$. Note the subscripts src and dest denote the propagation direction. If $tag_{rule}=c$, then
@@
$\frac{\partial tag_{n^{\prime}}}{\partial a_{n}}=\begin{cases}1,&\text{if }n^{\prime}=n\\ 0,&\text{if }n^{\prime}\neq n\end{cases}$ (8)
Then we focus on calculating gradients of $g_{e}$, which is more complicated. From Eq. 1, we know that $tag_{dest}^{new}$ is determined by $tag_{dest}$, $tag_{rule}$, and $g_{e}$. And $tag_{dest}$, $tag_{rule}$ are related to the propagation rates of previous events. Therefore, $tag_{dest}^{new}$ is influenced by the propagation rates of the current event (denoted by $g_{e}$) and all previous events (denoted by $g_{e^{\prime}}$) happened on $src$ and $dest$ nodes. Consequently, for each propagation, we must update $\partial tag_{dest}^{new}/\partial g_{e}$ and $\partial tag_{dest}^{new}/\partial g_{e^{\prime}}$ simultaneously. The case of $g_{e^{\prime}}$ is similar to that of $a_{n}$, we have
$\frac{\partial tag_{dest}^{new}}{\partial g_{e^{\prime}}}$ $=g_{e}\frac{\partial tag_{rule}}{\partial g_{e^{\prime}}}+(1-g_{e})\frac{\partial tag_{dest}}{\partial g_{e^{\prime}}}$ (9)
Calculating gradients of $g_{e}$ involves the Product Rule in calculus, which is
$\frac{\partial tag_{dest}^{new}}{\partial g_{e}}$ $=g_{e}\frac{\partial tag_{rule}}{\partial g_{e}}+(1-g_{e})\frac{\partial tag_{dest}}{\partial g_{e}}+tag_{rule}-tag_{dest}$ (10)
The full proof of Eq. 9 and Eq. 10 can be found in the Appendix §C.
@@
We use Eq. 11 to initialize the gradient of $g_{e}$ and update them according to Eq. 9 and  10.
In summary, to employ the gradient descent algorithm, we calculate the gradients of the adaptive parameters with respect to the loss. We first calculate the initial gradient for each node. Afterward, the gradients are updated according to our equations when the tags are propagated. Please note that unlike GNN models such as GCN [48] and GraphSage [49], which only consider the n-hop neighbor of a node, Captain fine-tunes every edge and node in the graph that can influence the detection results by propagating their gradients.
@@ IV-C3 Training and Testing
After calculating the gradients of the adaptive parameters with respect to the loss using the differentiable detection framework, we can now utilize the gradient descent algorithm to optimize the parameters. In machine learning, this process is referred to as “training”. The learned parameters are stored as a customized configuration, which is then used to set up Captain prior to testing.
As shown in Fig. 3, an epoch in training comprises the forward and backward propagation. Before training starts, all adaptive parameters, $A$, $G$, and $T$, are configured as the default settings $A_{0}$, $G_{0}$, and $T_{0}$. Instead of random initialization[50], we use the most conservative setting as the starting point to keep the sensitivity to the maliciousness during training.
When Captain processes the events, tags are propagated among the graph, updating the gradients according to Eq. 7,  9 and  10, and generate the detection results. Next, we calculate the loss and back-propagate the gradients of loss according to Eq. 5. Finally, parameters are updated in the opposite direction of the gradient as follows:
@@
where $p$ is the adaptive parameters (which could be $a_{n}$, $g_{e}$, or $thr_{e}$), $p_{new}$ is the new parameters, $p_{old}$ is the old one, $\frac{\partial Loss}{\partial p_{old}}$ is the gradient, and $l$ is the learning rate. We repeat the training process when the maximum epoch is reached or the changes in the result become sufficiently small. Before testing, we configure the parameters according to what we learned in the training stage. Then, Captain can process audit data and conduct detection as introduced in § IV-A.